X= 842, n= 1200, 95% CONFIDENCE


1
Expert's answer
2022-04-22T02:55:36-0400

"\hat p=\frac{842}{1200}=" 0.0702.

"95\%CI=(0.0702-1.96\sqrt{\frac{0.0702(1-0.0702)}{1200}},0.0702+1.96\sqrt{\frac{0.0702(1-0.0702)}{1200}})="

"=(0.6758,0.7276)."